# ChartBlender: An Interactive System for Authoring and Synchronizing Visualization Charts in Video

### v1

@@ 3 System Design and Overview
    Depth-Aware Embedding. Visualizations should be embedded into the video by aligning with camera depth and adopting a perspective consistent with the video scene. Such alignment ensures that the visualizations appear perceptually native to the scene, rather than as artificial post-production overlays.
@@
    Motion Synchronization. As the video plays, the visualizations should move in synchrony with the video, including the movement of both the camera and objects in the scene. To achieve this synchronization, it is necessary to track camera and object motion accurately.
@@ 4 Chart Embedder
In this section, we present a tracking algorithm for synchronizing visualizations with video motion. The algorithm supports two scenarios: (1) keeping visualizations stably anchored to the background as the camera moves (camera tracking), and (2) allowing visualizations to follow moving objects within the scene (object tracking).
As shown in Fig. 2, given an input video consisting of $N$ frames and a user-specified target point $(x,y)$, our goal is to generate a sequence of 3D pose matrices $O_{t}$. Each pose matrix represents the position and rotation of the tracked point in 3D space at frame $t$. It begins with depth estimation (Fig. 2 (a)) from the input video frames, which produces per-frame depth maps. These maps are used in two parallel modules: camera tracking (Fig. 2 (b)), which computes frame-to-frame transformations and accumulates them into a global camera pose; and object tracking (Fig. 2 (c)), which tracks a user-specified point across frames and combines its 2D trajectory with depth information to estimate a smoothed 3D object pose. Finally, the estimated object pose is applied to the visualization, which is then overlaid onto the video scene.
Figure 2: The pipeline of the Chart Embedder.
@@ 4.1 Depth Estimation
Our system employs a monocular depth estimation model, Metric3D[38], to generate per-frame metric depth maps directly from original video frames. The major advantages of adopting Metric3D are twofold: (1) it predicts absolute metric depth without requiring ground-truth depth or camera intrinsics, enabling zero-shot deployment on in-the-wild videos; (2) it achieves high generalization capability across scenes, owing to its large-scale training on diverse data.
@@
Given these benefits, we apply Metric3D to each of the $N$ video frames to generate a sequence of dense depth maps $D_{n}$, $n\in{0,1...,N-1}$. For a user-specified 2D point $(x,y)$, we can extract its first frame’s depth value from the depth map as $d=D_{0}(x,y)$. This depth information serves as a crucial foundation for the subsequent camera tracking and object tracking processes, enabling accurate spatial reasoning in both modules.
4.2 Camera Tracking
In ChartBlender, camera tracking ensures that embedded 3D visualizations remain spatially consistent with the static element throughout the video (Fig. 3(a)). To achieve this, we estimate the camera’s motion through a three-step process. First, we perform (1) relative pose estimation to compute the geometric transformation between adjacent video frames, capturing how the camera moves over time. This information is then used in (2) camera pose estimation, accumulating relative transformations to recover the camera’s trajectory in the scene. Finally, we conduct (3) 3D pose recovery to infer the camera’s position and orientation in 3D space, enabling visualizations to be accurately projected.
@@ 4.2 Camera Tracking
Relative pose estimation. To understand how the camera moves from frame to frame, we should estimate the relative transformation between each pair of consecutive video frames. These pairwise transformations form the foundation for reconstructing the full camera trajectory. We employ an RGBD odometry model provided by Open3D [39, 40] to compute the relative camera motion between adjacent RGBD frames. Given two consecutive frames $n$ and $n+1$, RGBD odometry estimates a transformation matrix. Specifically, we extract a transformation matrix $T_{n\rightarrow n+1}$ that aligns the RGB and depth images. Each transformation is a $4\times 4$ homogeneous transformation matrix composed of a rotation matrix $R_{n\rightarrow n+1}$ and a translation vector $t_{n\rightarrow n+1}$:
$T_{n\rightarrow n+1}=\left[\begin{array}[]{cc}R_{n\rightarrow n+1}&t_{n\rightarrow n+1}\\ 0&1\end{array}\right]$ (1)
@@
Camera pose estimation. While relative poses describe local motion between frames, embedding visualizations in a stable, world-aligned coordinate system requires each frame’s absolute camera pose. This enables consistent rendering of the 3D object across the video sequence. We compute the absolute camera pose $M_{n}$ at each frame $n$ by chaining all previous relative transformations starting from the first frame:
$M_{n}=T_{1\rightarrow 2}T_{2\rightarrow 3}\cdots T_{{n-1}\rightarrow n}$ (2)
This accumulation step reconstructs the global camera trajectory, allowing us to express each frame’s viewpoint in the same coordinate system.
@@
3D pose recovery. To render the 3D visualization as if it were physically anchored in the scene, we must compute its position in the global world coordinate system. Given a user-specified 2D point $(x,y)$ in the image and its corresponding predicted depth $d$, we first back-project the point into the 3D camera coordinate system using the inverse of the camera intrinsic matrix $K^{-1}$. The resulting 3D point in the camera frame is:
$P_{\text{n}}=K^{-1}\cdot\left(d_{0}\cdot\tilde{o}_{n}\right)$ (3)
where $P_{n}$ denotes the 3D position of a tracked point at frame $n$, $\tilde{o}_{n}=[x,y,1]^{T}$ is the homogeneous pixel coordinate. Next, we transform this point into the world coordinate system using the inverse of the absolute camera pose $M_{n}$, computed from accumulated relative poses. The object’s global pose $O_{n}$ is then given by:
$O_{n}=M_{n}^{-1}\cdot P_{\text{n}}$ (4)
This transformation places the 3D visualization at a consistent location in the world coordinate system, regardless of camera motion. It ensures that the visualization remains spatially stable and visually aligned with the static background in the video.
4.3 Object Tracking
In ChartBlender, the goal of object tracking is to enable visualizations to remain spatially aligned with moving elements in the video scene (Fig. 3(b)). To achieve this, we estimate the 3D motion of target objects from monocular video input through a three-stage pipeline. First, we perform (1) 2D point tracking to extract the object’s pixel-wise trajectory across frames, capturing how it moves within the image plane. This trajectory serves as the basis for the second stage, (2) 3D translation estimation, which lifts the 2D motion into 3D space to approximate the object’s position relative to the camera over time, allowing visualizations to maintain consistent placement as the object moves. Finally, we perform (3) 3D rotation estimation to recover the object’s orientation changes, ensuring that embedded visualizations can rotate coherently with the object’s local motion, further enhancing the perceptual integration.
@@ 4.3 Object Tracking
2D point tracking. We begin by identifying salient points on dynamic objects and tracking their 2D positions across frames. To accomplish this, we use BootsTAP [41], a high-precision keypoint tracking algorithm that leverages dense visual correspondence estimation. For each target point $(x,y)$, the tracker generates a pixel-level trajectory over time, denoted as $p_{n}=(u_{n},v_{n})\rightarrow p_{n+1}=(u_{n+1},v_{n+1})$. The 2D trajectories capture the apparent motion of object features in image space and form the foundation for estimating real-world displacement in 3D, which is critical for subsequent motion reconstruction.
3D translation estimation. To reconstruct the spatial motion of tracked points, we lift the 2D trajectories to 3D space. For each frame, the 2D trajectory is projected into 3D space using the depth map of the tracked point in the camera coordinate system.
$d_{n}=D_{n}(p_{n})$ (5)
However, monocular depth estimates are inherently noisy and temporally unstable due to factors such as low-texture regions, occlusions, and variations in lighting. These inconsistencies can lead to jitter and discontinuities in the reconstructed 3D trajectories, especially when tracking fine-grained object motion. Such artifacts degrade both the physical plausibility and visual coherence of the motion path, undermining downstream tasks like 3D alignment and visualization anchoring. To address this, we apply temporal smoothing to the depth signal using quadratic regression over a sliding window:
$\hat{d_{n}}=\text{QuadraticRegression}(d_{n})$ (6)
As previously defined in Equation (3), we recover the 3D translation vector $t_{n}$ of the target point in the camera coordinate system by back-projecting its 2D image coordinate using the estimated depth and the intrinsic matrix K.
@@
3D rotation estimation. Finally, we estimate object rotation based on the direction of motion in 3D space to ensure that visualizations remain properly oriented when attached to moving objects. Specifically, we define the per-frame motion vector as $V_{n}=P_{n+1}-P_{n}$. We obtain the rotation vector $R_{n}$ by constraining the object’s local X axis to follow $V_{n}$. To support stable rotation estimation, we first regularize the raw 3D point trajectories by minimizing sudden changes in displacement. This is implemented as a second-order temporal smoothing loss:
@@
Next, we apply the same principle in the velocity domain to ensure that the direction of motion also varies smoothly over time. The corresponding velocity regularization is formulated as:
$\mathcal{L}(V_{n})=\sum_{n}\|V_{n+1}-V_{n}\|^{2}+\lambda\sum_{t}\|V_{n+1}-2V_{n}+V_{n-1}\|^{2}$ (8)
This second-order smoothing constraint encourages both consistent motion direction and smooth acceleration profiles, mitigating jitter and improving temporal coherence in the estimated trajectories. Given the smoothed translation vector and rotation vector, the final object pose at each frame is defined as $O_{n}=\{R_{n}\mid t_{n}\}$.
@@ 5.2 Visualization Design
Our goal is to design visualizations that are well-suited for integration within videos. This requires rethinking how charts are structured and styled to maintain both visual clarity and contextual harmony in videos[42]. To this end, we collected a total of 200 high-quality videos and manually segmented them into 502 clips, each containing at least one embedded visualization. As shown in Fig. 6, these videos span ten diverse domains. We also analyzed the most commonly used chart types across the dataset. Furthermore, we identified several recurring design patterns in how visualizations are embedded within real-world video content. These patterns are summarized as follows.
Figure 5: Design space for visualizations suitable for embedding in video, which illustrates five representative visualization forms combined with two motion anchoring strategies (object tracking and camera tracking). It reveals common design patterns for how visualizations are structured and integrated within dynamic video scenes across diverse real-world applications.
@@
First, a majority of charts (130 cases) incorporate background to improve readability, with a subset (21) employing semi-transparent backgrounds to maintain visual access to the underlying scene. Notably, gridlines were selectively applied to support data interpretation, with horizontal guides appearing more frequently (109) than vertical ones (43), while 64 charts omitted them entirely. Second, very few charts (only 4) included legends, indicating a general preference for minimalist presentations in time-constrained video contexts. Nevertheless, essential structural components such as axes (retained in 171 charts) and titles (present in 168 charts) were commonly preserved. Third, when visualizations were associated with specific objects in the scene, designers frequently employed visual linking strategies to establish a clear connection between the chart and its target (68). These strategies included directional cues such as arrows, lines, and bounding boxes. The choice of pointer type often varied depending on scene complexity: in simple scenes with few moving objects, minimal indicators such as triangles were typically sufficient; in more complex scenes with multiple overlapping motions, designers commonly used bounding boxes or anchored pointers to more clearly distinguish the tracked target.
These findings highlight a set of implicit yet widely adopted design patterns for video-suited visualizations design, which we synthesize into a structured design space, as illustrated in Fig. 5. This design space is organized along two key dimensions: the motion mode and the visualization type. It distinguishes between two modes of motion (object tracking and camera tracking) corresponding to whether the visualization follows a moving target in the scene or remains fixed relative to the global environment. Horizontally, it categorizes common visualization types, including text annotations, line charts, area charts, bar charts, and 3D charts, as well as their corresponding linking and layout strategies (e.g., aiming lines, bounding boxes, or indicators). Based on this design space, we developed a visualization chart library using Three.js, providing a collection of chart templates optimized for embedding in dynamic video scenes.
@@ 5.3 Interaction
To support efficient visualization authoring and seamless synchronization within video, we propose a set of interactions in ChartBlender:
Upload. Users can upload datasets and video files either by clicking the upload button or by dragging and dropping files into the designated panel.
Data Mapping. Data attributes can be mapped to visual channels through direct manipulation: users drag data fields onto chart elements to define encoding relationships.
Zoom and Pan. Users can interactively adjust the visualization view: left-click and drag to reposition charts, scroll to zoom in and out, and right-click and drag to rotate.
Chart and Style Switching. Users can switch between different chart types via tabs in the interface. They can also adjust visual styling, such as color schemes, by clicking style options or using a color picker tool.
Anchor Selection. To ensure accurate chart placement, users can click on static or dynamic objects in the video to define visual anchor points. These anchors are used to bind visualizations to specific elements in the scene, supporting consistent spatial alignment across frames.
Track Editing. Users can manage chart tracks via the timeline view. Tracks can be added or removed by clicking the corresponding buttons in the toolbar on the left. On the right, users can adjust the timing and duration of each chart by dragging chart tracks directly along the timeline.

### v2

@@ 3.2 Design Requirements
    Depth-Aware Embedding. Visualizations should be embedded into the video by aligning with camera depth and adopting a perspective consistent with the video scene. Such alignment ensures that the visualizations appear seamless to the scene, rather than as artificial post-production overlays.
@@
    Motion Synchronization. As the video plays, the visualizations should move in synchrony with the video, including the movement of both the camera and objects in the scene. To achieve this synchronization, it is necessary to accurately track both camera and object motion.
@@ 4 Chart Embedder
The Chart Embedder module supports two tracking pipelines: camera tracking and object tracking. Camera tracking keeps an embedded 3D visualization spatially consistent with a static scene element throughout the video (Fig. 3(a)), whereas object tracking keeps the visualization aligned with a moving target (Fig. 3(b)).
As shown in Fig. 4, both pipelines start with (A) depth estimation, which provides per-frame depth cues needed to lift 2D image measurements into 3D and to render charts with perspective-consistent projection. They then differ in their motion estimation stage: (B1) camera pose estimation for camera tracking recovers the camera trajectory so a chart can remain anchored to a fixed location in the scene despite camera motion, whereas (B2) object trajectory estimation for object tracking recovers the 3D motion of a target so a chart can follow the object over time.
Figure 4: The two pipelines of the Chart Embedder. The camera tracking pipeline consists of (A) depth estimation followed by (B1) camera pose estimation. The object tracking pipeline consists of (A) depth estimation followed by (B2) object trajectory estimation.
@@ 4.1 Depth Estimation
Unlike general-purpose 2D trackers that return image-plane trajectories, chart embedding requires 3D, scene-referenced motion signals: we need depth to lift user-selected points from pixels into 3D space and to compute poses that keep a chart perspective-consistent and spatially stable under camera motion. To obtain the required spatial context, our system uses a monocular depth estimation model, Metric3D [yin2023metric], to predict metric depth maps for each video frame. We adopt Metric3D for two reasons: (1) it predicts absolute metric depth without requiring ground-truth depth or camera intrinsics, enabling zero-shot deployment on in-the-wild videos; and (2) it generalizes well across diverse scenes due to large-scale training data.
@@
Given these benefits, we apply Metric3D to each of the $N$ frames to obtain a sequence of dense depth maps ${D_{n}}_{n=0}^{N-1}$. For a user-specified 2D point $(x,y)$, we extract the depth value in the first frame as $d=D_{0}(x,y)$. This depth signal provides the foundation for subsequent camera and object tracking, enabling spatial reasoning in both modules.
4.2 Camera Pose Estimation
In ChartBlender, camera pose estimation aims to recover a consistent camera trajectory across frames so that a visualization anchored to a static scene location remains stable under camera motion. We decompose this problem into three steps. First, we perform (1) relative pose estimation to compute the geometric transformation between adjacent frames, capturing local camera motion that can be robustly inferred from short temporal baselines. Second, we perform (2) global pose estimation by accumulating these relative transformations to obtain a per-frame camera pose in a shared coordinate system, which is required for world-consistent anchoring over time. Finally, we perform (3) 3D pose recovery to lift the user-specified 2D anchor (with predicted depth) into a world-space reference and derive the corresponding per-frame chart pose, enabling accurate 3D placement and projection throughout the video.
@@ 4.2 Camera Pose Estimation
Relative pose estimation. To understand how the camera moves from frame to frame, we should estimate the relative transformation between each pair of consecutive video frames. These pairwise transformations form the foundation for reconstructing the full camera trajectory. We employ an RGBD odometry model provided by Open3D [Zhou2018, park2017colored] to compute the relative camera motion between adjacent RGBD frames. Given two consecutive frames $n$ and $n+1$, RGBD odometry estimates a transformation matrix. Specifically, we extract a transformation matrix $T_{n\rightarrow n+1}$ that aligns the RGB and depth images. Each transformation composed of a rotation matrix $R_{n\rightarrow n+1}$ and a translation vector $t_{n\rightarrow n+1}$.
@@
Global pose estimation. While relative poses describe local motion between frames, embedding visualizations in a stable, world-aligned coordinate system requires each frame’s absolute camera pose. This enables a consistent rendering of the 3D object throughout the video sequence. We compute the absolute camera pose $M_{n}$ at each frame $n$ by chaining all previous relative transformations starting from the first frame: $M_{n}=T_{1\rightarrow 2}T_{2\rightarrow 3}\cdots T_{{n-1}\rightarrow n}$ This accumulation step reconstructs the global camera trajectory, allowing us to express each frame’s viewpoint in the same coordinate system.
@@
3D pose recovery. To render the 3D visualization as if it were physically anchored in the scene, we should compute its position in the global world coordinate system. Given a user-specified 2D point $(x,y)$ in the image and its predicted corresponding depth $d$ predicted by the depth estimation module, we first reproject the point back into the 3D camera coordinate system using the inverse of the intrinsic matrix of the camera $K^{-1}$. Next, we transform this point into the world coordinate system using the inverse of the absolute camera pose $M_{n}$, computed from accumulated relative poses. This transformation places the 3D visualization at a consistent location in the world coordinate system, regardless of camera motion. It ensures that the visualization remains spatially stable and visually aligned with the static background in the video.
4.3 Object Trajectory Estimation
Object trajectory estimation aims to recover a per-frame 3D pose for a moving target so that an attached visualization follows the object with stable position and coherent orientation over time. We decompose this problem into three stages. Specifically, we estimate per-frame object poses via (1) 2D point tracking, which extracts a pixel trajectory; (2) 3D translation estimation, which combines the tracked 2D positions with the predicted depth to recover the object’s 3D displacement; and (3) 3D rotation estimation, which infers a smooth orientation change to keep the visualization consistently aligned with the object’s motion. Together, these stages produce per-frame poses that drive motion-synchronized chart embedding for dynamic targets.
@@ 4.3 Object Trajectory Estimation
2D point tracking. We identify salient points on the target object and track their 2D locations across frames. We use BootsTAP [doersch2024bootstap], a high-precision keypoint tracker based on dense visual correspondence. For a target point $(x,y)$, the tracker outputs a pixel-level trajectory $p_{n}=(u_{n},v_{n})$ across frames. These trajectories capture apparent motion in image space and serve as the basis for recovering 3D displacement.
3D translation estimation. We lift 2D trajectories into 3D by combining them with per-frame depth. For each frame, we obtain the depth at the tracked pixel from the corresponding depth map, $d_{n}=D_{n}(p_{n})$, and back-project $p_{n}$ to a 3D point $P_{n}$ in camera coordinates. Because monocular depth can be noisy and temporally inconsistent (e.g., due to low-texture regions, occlusion, or lighting changes), the recovered 3D trajectory may exhibit jitter. We therefore smooth the depth signal over time using quadratic regression within a sliding window, and then compute the per-frame translation vector $t_{n}$ via back-projection with the smoothed depth and intrinsic matrix $K$.
@@
3D rotation estimation. We estimate object rotation from the direction of motion in 3D so that the visualization remains properly oriented as the object moves. We define the per-frame motion vector as $V_{n}=P_{n+1}-P_{n}$, and obtain a rotation $R_{n}$ by constraining the object’s local $x$-axis to align with $V_{n}$. To stabilize rotation, we first regularize 3D point trajectories by minimizing sudden changes in displacement using a second-order temporal smoothing objective:
@@
where $\lambda$ controls the strength of the acceleration penalty. We apply the same principle in the velocity domain to encourage smooth motion direction over time, mitigating jitter and improving temporal coherence. Given the estimated translation $t_{n}$ and rotation $R_{n}$, the per-frame object pose is defined as $O_{n}=\{R_{n}\mid t_{n}\}$.
@@ 5.2 Visualization Design
We observed that visualizations appearing in videos often differ from conventional charts. To understand how embedded visualizations are designed in practice and to inform the design of our chart library, we followed a systematic workflow: we collected a real-world video corpus and synthesized recurring strategies into a design space of embedded-visualization principles, offering practical implications for both researchers and designers.
5.2.1 Methods
To ground our design in practice, we assembled a corpus of 200 publicly available, high-quality videos featuring embedded visualizations from online platforms (e.g., YouTube and Bilibili) and official outlets in domains such as news, sports, technology, and education. Each video was manually segmented at visualization entry and exit points, yielding 502 clips containing at least one embedded visualization. Two trained annotators coded each clip using a codebook that combined deductive categories (chart type, 2D/3D representation, spatial relation to the scene, motion coupling to camera/object, occlusion handling, and labeling/styling) with inductive categories for emergent patterns. After an initial calibration phase, disagreements were resolved through discussion. We also recorded domain metadata spanning ten categories (e.g., economy, sports, ecology, and transportation), as summarized in Fig. 7.
Figure 6: Design space for visualizations suitable for embedding in video, which illustrates five representative visualization forms combined with two motion anchoring strategies. It reveals common design patterns for how visualizations are structured and integrated within dynamic video scenes across diverse real-world applications.
@@ 5.2.1 Methods
Furthermore, we identified several recurring design patterns in how visualizations are embedded within real-world video content. These patterns are summarized as follows. (1) Legibility. Background panels are widely adopted to enhance readability (130/502), with a minority employing semi-transparent panels to maintain visual context (21/502). Gridlines are used selectively: horizontal guides appear more frequently (109) than vertical ones (43), while a notable number of charts omit gridlines entirely (64). (2) Referential Linking. When visualizations correspond to specific objects in the scene, designers often employ visual linking techniques (e.g., arrows, lines, or bounding boxes) to establish clear referential ties (68). The choice of pointer type tends to adapt to scene complexity: in simple scenes with sparse motion, minimal markers like triangles often suffice; in cluttered environments with overlapping movements, bounding boxes or anchored pointers are preferred to unambiguously identify the target. (3) Simplicity. Legends are rarely used (only 4 instances), reflecting a preference for minimalism in video formats. In contrast, essential structural elements such as axes (retained in 171 charts) and titles (present in 168 charts) are commonly preserved to maintain baseline interpretability.
5.2.2 Design Space
These findings highlight a set of implicit yet widely adopted design patterns for video-suited visualizations design, which we synthesize into a structured design space, as illustrated in Fig. 6. Notably, although we identified a broad spectrum of ten chart types across diverse application scenarios (as shown in Fig. 7), we focused the design space on the five most commonly used chart types in the corpus to maximize generalizability.
The space is organized along three primary dimensions: (i) chart type, covering five common forms, including data, text, line charts, bar charts, and area charts; (ii) motion mode, including object-tracking and camera-tracking embeddings; and (iii) temporal behavior shown at the bottom of each schematic. We describe each visual element by its on-screen duration and animation scheme.
Importantly, these dimensions are not independent; instead, they co-vary with production constraints and communicative intent in systematic ways. First, the choice of motion mode usually reflects what the visualization refers to. When the message is about a specific entity in the scene (e.g., a person, a vehicle, or a building), creators tend to use object-tracking so the overlay stays visually attached to the target. In these cases, the visualization is often kept compact (short text/data, small charts) and paired with explicit cues such as arrows or boxes. In contrast, when the visualization summarizes the scene or provides background context (e.g., overall trends or comparisons), creators more often use camera-tracking overlays. This allows larger chart panels and reduces the risk that the overlay jitters or drifts with local object motion. Second, temporal behavior is largely driven by attention and editing rhythm. Embedded charts rarely stay static on screen. Instead, they appear when the narration or shot content calls for data, remain long enough to be read, and then disappear to return attention to the footage. As a result, the most common animations are subtle and functional (e.g., fade-in/out or simple unveil), because aggressive animation competes with motion in the video and makes reading harder.
Finally, these observations explain why a design space is useful for system building. In professional workflows, creators repeatedly adjust the same small set of things: what the overlay is anchored to, how strictly it tracks, where it sits to avoid covering important content, how strong the background panel should be for legibility, and when it enters/exits. The design space guided our system’s chart library design by turning the most frequently adjusted parameters into explicit controls and offering presets for common embedded chart styles.
@@ 5.3 Interaction
To support efficient authoring and reliable synchronization, ChartBlender provides an integrated workflow that covers data binding, spatial anchoring, and temporal configuration. Key interaction techniques include:
Upload and Data Mapping. Users upload video footage and datasets via drag-and-drop or file selection. When timestamps are available, the system aligns data records with the video timeline. Users map data fields to visual channels by dragging them onto chart components, and the preview updates immediately to reflect mapping changes.
One-click Tracking Anchor. Users specify a tracking anchor by simply clicking a pixel in the video frame. Based on the selected motion mode, the system either (1) attaches the visualization to a moving target via Object Tracking or (2) keeps the chart anchored to the static scene via Camera Tracking in world coordinates.
3D Adjustment. Users pan, zoom, and rotate the chart in 3D to fine-tune placement: dragging pans, scrolling zooms, and right-dragging rotates the view. These operations work together with anchoring to achieve stable alignment and perspective-consistent embedding.
Real-time Preview. ChartBlender provides continuous, in-context preview throughout authoring. While editing, users can immediately see changes to chart styling and 3D viewpoint directly on the video. After rendering, users can play back the result to preview the chart’s motion in real time and verify alignment and visual coherence over time.
Multi-chart Editing. Users manage when each visualization appears using the timeline. Multiple charts are organized as separate tracks and can be added or deleted with simple timeline operations. Users adjust a chart segment’s entry, exit, and duration by dragging its endpoints, and selecting a segment brings that chart into focus for independent editing of its data mapping, style, and placement.